OpenStudy (anonymous):

Find 4 consecutive even integers where the product of the two smaller numbers is 56 less than the product of the two larger numbers. @tcarroll010 @matty546

OpenStudy (anonymous):

(2x)(2x + 2) + 56 = (2x + 4)(2x + 6) 4x^2 + 4x + 56 = 4x^2 + 20x + 24 16x = 32 x = 2 So, since your first number is 2x: 4, 6, 8, 10

OpenStudy (anonymous):

You have to use 2x, 2x + 2, 2x + 4, and 2x + 6 not x, x + 1, x + 2, x + 3 because you have to force "evenness". That's why you need everything multiplied by "2".

OpenStudy (anonymous):

And since this works out to be: 4, 6, 8, and 10 you can verify this by: (4 x 6) + 56 = (8 x 10) So, it works.

OpenStudy (anonymous):

All good now, @shortie212 ?

OpenStudy (anonymous):

yea i am gann posted more @tcarroll010